# F-91191 Gif-sur-Yvette cedex, France

### v1

@@ 3 The asymptotic expansion of expectation values
While Eq. (22) defines the metric sourced by a scattering process at all space-time points, carrying out the integrals can be difficult. For our purpose, however, such a calculation is unnecessary. Indeed, a large-$r$ asymptotic expansion of this metric should be sufficient to probe its peeling properties. Furthermore, it is more convenient to evaluate Newman-Penrose scalars to probe gauge-invariant properties of spacetime. In this section, we begin by discussing the asymptotic expansion of Eq. (22) large values of $r$. Then, we upgrade the analysis to cover the evaluation of the asymptotic expansion of derivative operators acting on the metric and, in particular, of the Newman-Penrose scalars.
@@
$k^{\mu}=\omega\,t^{\mu}+\mathsf{k}\,n^{\mu}\ ,\quad\text{where}\ n^{2}=-1$ (29)
@@
$\begin{split}\int\!\frac{d^{d}k}{(2\pi)^{d}}\frac{e^{-ikx}\,{\widetilde{h}}^{\mu\nu}(k)}{(k_{0}+i\epsilon)^{2}-\vec{k}^{2}}=\int_{-\infty}^{+\infty}\!\frac{d\omega}{2\pi}\int_{0}^{\infty}\!\frac{d\mathsf{k}}{2\pi}\frac{e^{-i\omega t\cdot x}\mathsf{k}^{d-2}}{(\omega+i\epsilon)^{2}-\mathsf{k}^{2}}\int\!dn\,e^{-i\mathsf{k}n\cdot x}{\widetilde{h}}^{\mu\nu}(k)\ ,\end{split}$ (30)
@@
The contour for the $\omega$ integral, as defined by the $i\epsilon$ prescription in Eq. (22), encodes our desired (retarded) boundary conditions. The integration can be carried out via a contour deformation. In particular, by deforming the $\omega$ integration contour to the lower-half plane, we can pick up contributions from the singularities of the source and the external (retarded) propagator. This integral can typically be performed exactly.
@@
Using this additional direction, the graviton momentum can be parametrized as
@@
where $k^{\mu}$ in the argument of the current is given by Eq. (32), and $dn_{\perp}$ is the measure on the corresponding (d3)(d-3)-dimensional sphere. The $z$ angular integration can be carried out either together with or after the $\omega$ integration.
The $\mathsf{k}$ integration is more subtle. At large distances, the integrand depends on two parametrically-separated scales: that of the system, ${\cal O}(b)\sim{\cal O}(R_{i})$, and that of the observer, ${\cal O}(|\vec{x}|)$. Depending on the scaling of $\mathsf{k}$ with the observer scale, the contribution of the oscillatory factor is different. Thus, we have two distinct integration regions,
@@
Since the complete $\mathsf{k}$ integral is convergent both at small $\mathsf{k}$ (because of the known soft behavior of ${\widetilde{h}}^{\mu\nu}(k)$) and at large $\mathsf{k}$ (because of the exponential suppression ensuing from the typical frequency of ${\widetilde{h}}^{\mu\nu}(k)$ being ${\cal O}(1/b)$), a divergence in one of the two regions should be regularized by the cutoff $\lambda$ and must cancel against a similar divergence in the other region. Both of these regions contribute to the asymptotic expansion at large $r$ and need to be considered separately. From the perspective of the method of regions in dimensional regularization Beneke:1997zp ; Jantzen:2011nz , the graviton momentum scales at large $|\vec{x}|$ in these regions are
$\text{Radiation region}:\mathsf{k}^{\mu}\sim{\cal O}(|\vec{x}|^{0})\qquad\qquad\text{Coulomb region}:\mathsf{k}^{\mu}\sim{\cal O}(|\vec{x}|^{-1}).$ (35)
@@ 3.1 The radiation region
In this region, it is convenient to deform the $z$ contour into two contours parallel to the imaginary axis and connected at infinity, as in Fig. 2. The result of the integration is
@@
which is valid for a vector $n$ parametrizing the unit $n$-sphere. In our case, $n_{\perp}$ is further constrained to be orthogonal on $t^{\mu}$ and $r^{\mu}$; relevant projector enforcing this transversality is
@@
and the integral is a function of only $a^{\mu}_{\perp}=\eta_{\perp}^{\mu\nu}a_{\nu}$. The integrals over $\mathsf{k}$ and $\omega$ are more interesting. They depend on the details of the source ${\widetilde{h}}^{\mu\nu}(k)\sim J^{\mu\nu}(k)$ and are evaluated on a case-by-case basis. At tree-level, the $\omega$ integral can be evaluated via the residue theorem. Beyond tree level, the integral over $\omega$ is complicated by the appearance of logarithmic branch-points at $\omega=\mathsf{k}$ in $J^{\mu\nu}(k)$ in $d=4$. We leave the evaluation of these integrals to Sections 4.3 and 4.4. As a cross-check, we will also show there that the leading contribution in $|\vec{x}|^{-1}$ reproduces the waveform at future null infinity, originally obtained in Ref. Cristofoli:2021vyo via a saddle-point approximation.
@@ 3.2 The Coulomb region
In this region, the external graviton momentum is taken to be parametrically soft, ${\cal O}(|\vec{x}|^{-1})$; moreover, since we are interested in the metric at large distance towards future null infinity, we will assume that, while $\omega\neq\mathsf{k}$, they are of the same order, $\omega\sim\mathsf{k}$. Under these assumptions the source ${\widetilde{h}}^{\mu\nu}(k)$ can be expanded in powers of $\omega$ and $\mathsf{k}$. Displaying only the $\mathsf{k}$ dependence, ${\widetilde{h}}^{\mu\nu}(k)$ schematically has the general form Georgoudis:2023eke
@@
Since $\omega,\mathsf{k}\sim|\vec{x}|^{-1}$ in the Coulomb region, the phase of the Fourier transform in Eq. (3.1) can be further simplified by expanding it in the small quantity $\omega u$:
@@
Subleading orders in this expansion contribute beyond the leading order in the $1/|\vec{x}|$ expansion. Following the usual procedure provided by the method of regions in dimensional regularisation, we take the upper limit of the integration domain to $\lambda\rightarrow\infty$ after series expanding the integrand.
@@
In the following, we will focus only on the leading and subleading soft contributions, corresponding to the first term in each of the two sums in Eq. (39). We will study both terms at all orders in the coupling. We find that only the leading orders in the soft expansion contribute to peeling violation. In particular, we will see that the leading $\mathsf{k}^{-1}$ contribution satisfies non-trivially the peeling conditions (3) and that the leading term in the ultra-soft exchange region reproduces at leading order in $G$ the peeling violation (116), identified in Ref. Christodoulou:2002 .
@@
are chosen so that the integrals in $J^{\mu\nu}(q_{1},q_{2})$ exhibit a clean separation of the classical orders. $J^{\mu\nu}(q_{1},q_{2})$ is the final-state graviton 1-point function in momentum space, and the two regions correspond to the following rescalings Georgoudis:2023eke :
$\text{SE}:~(k^{\mu},q_{i}^{\mu})\sim(|\vec{x}|^{-1},1)\qquad\qquad\text{U-SE}:~(k^{\mu},q_{i}^{\mu})\sim(|\vec{x}|^{-1},|\vec{x}|^{-1})$ (43)
expanded at large $x$. SE and U-SE stand for soft-exchange and ultra-soft-exchange, respectively. As usual in the method of regions in dimensional regularization, the integrals are evaluated by expanding at small $\omega$ in each region and integrating over the entire integration domain.
@@ 3.3 Ansatz for metric
As discussed, gauge invariance of the Einstein-Hilbert action coupled to the source $J^{\mu\nu}$ requires that the source be transverse. We thus parametrize it in terms of manifestly-transverse tensors. At lowest order in the graviton momentum, there are six such tensors,
@@
The coefficients $\alpha_{i}$ turn out to be finite throughout the radiation region, $k^{\mu}\sim{\cal O}(|\vec{x}|^{0})$, a property that we will leverage in later sections. To see this, recall that singularities of amplitudes (and of Green’s functions) correspond to physical processes and, as such, can be understood using a combination of unitarity arguments and analytic continuation DeAngelis:2023lvf ; Brunello:2024ibk . In two-body scattering, the relevant invariants appearing in the impact-parameter-space form of the amplitude are $u_{1,2}\cdot k$, $b\cdot k$, $b^{2}$, $u_{1}\cdot u_{2}$, and $k^{2}$ (with the latter vanishing on-shell). The large distance expansion in the radiation region, corresponds to an essentially null momentum $k^{\mu}$, aligned with the null direction connecting the scattering event to the observer. In this limit, no singularities can arise. Indeed, the $u_{i}^{\mu}$ are time-like so divergences in $u_{i}\cdot k$ reduce to the standard soft singularities in $\omega$. Furthermore, no singularities can occur in $k\cdot b$, as the observer may lie anywhere on the celestial sphere and the impact parameter remains arbitrary. Moreover, since none of the tensor structures in Eq. (46) vanish as $z\rightarrow 1$, it follows that the coefficient functions $\alpha_{i}$ remain finite in this limit. These expectations can be confirmed explicitly from the known LO and NLO post-Minkowskian scattering waveforms Kovacs:1977uw ; Jakobsen:2021smu ; Herderschee:2023fxh ; Brandhuber:2023hhy ; Georgoudis:2023lgf ; Bini:2023fiz .
@@ 4 The peeling property by region
Using the methods outlined in the previous sections, we explicitly evaluate the asymptotic expansion of the Fourier transform (30). We discuss separately the Coulomb and radiation regions, focusing in Secs. 4.1, 4.2.1 and 4.3 on terms which are at rational in $\mathsf{k}$ and $\omega$. These originate from the soft-exchange part (43) of the Coulomb region (35), and from the terms which are analytic in $k^{2}$ in the radiation region (35). We will prove that such terms satisfy Sachs’s peeling property (3) as a consequence of nontrivial cancellations. We then discuss the remaining contributions – the ultra-soft-exchange part (43) of the Coulomb region (35) in Sec. 4.2.2 and the non-analytic terms in $k^{2}$ in the radiation region (35) in Sec. 4.4. The former yields violation of the peeling property already identified at leading order in Newton’s constantin Refs. Christodoulou:2002 . The latter, which we discuss in Sec. 4.4, introduce a new stronger violation.
@@ 4.1 Asymptotic behavior of the waveform: the gravitational memory
is on the one hand subleading in the soft limit and, on the other, it is analytic. These properties allow us to ignore the difference at leading order. Direct power counting indicated that the off-shell deformation contributes only suppressed terms in the large-$|\vec{x}|$ expansion; moreover, as we will see, the fact that they are analytic will also allow us to ignore them for the study of peeling properties. The leading soft behavior of the amplitude captures the information about asymptotic large retarded times $|u|$, i.e. the so-called memory effect.
The soft theorem for the two-body-scattering in-in correlator is [^13] Such a soft theorem for the two-body scattering can be inferred e.g from the result of Ref. Saha:2019tub : the leading soft theorem is given by $J^{\mu\nu}(k)=i\frac{\kappa^{2}}{2}\sum_{i=1}^{n+m}\frac{p_{i}^{\mu}p_{i}^{\nu}}{p_{i}\cdot k+i\epsilon}\ ,$ (54) where $n$ and $m$ are the number of incoming and outgoing particles, respectively (either massive or massless; for the incoming particles the $p_{i}$ are intended as minus their momenta). As our focus on the two-body scattering, we have to sum over the momenta of the incoming and outgoing massive states. During the scattering process, an certain amount of radiation is emitted and its total momentum is fixed by momentum conservation. Requiring gauge invariance to hold in the soft limit, fixes the fifth term to be of the form in Eq. (55). Note also that we work in a mostly-minus metric convention while Ref. Saha:2019tub has the opposite, mostly-plus metric convention.
$J^{\mu\nu}(k)$ $=i\frac{\kappa^{2}}{2}\left[-\frac{p_{1}^{\mu}p_{1}^{\nu}}{p_{1}\cdot k-i\epsilon}-\frac{p_{2}^{\mu}p_{2}^{\nu}}{p_{2}\cdot k-i\epsilon}+\frac{p_{1}^{\prime\mu}p_{1}^{\prime\nu}}{p_{1}^{\prime}\cdot k+i\epsilon}+\frac{p_{2}^{\prime\mu}p_{2}^{\prime\nu}}{p_{2}^{\prime}\cdot k+i\epsilon}+\frac{K^{\mu}K^{\nu}}{K\cdot k+i\epsilon}\right]$ (55)
$+\mathcal{O}(\mathsf{k}^{0},\mathsf{k}^{d-4})\ ,$ (56)
where $p_{1,2}^{\mu}$ are the two incoming matter momenta, $p_{1,2}^{\prime\mu}$ are their (exact) outgoing momenta, and $K^{\mu}$ is the four-momentum carried by the emitted radiation during the scattering process. The outgoing matter momenta are determined by the details of the interactions between the two bodies, and we will not need their precise form, which is known only perturbatively. For example, at leading order in Newton’s constant
@@
$p_{1}^{\mu}+p_{2}^{\mu}=p_{1}^{\prime\mu}+p_{2}^{\prime\mu}+K^{\mu}\ .$ (58)
It is convenient to introduce the shorthand notation $Q_{i}^{\mu}$’s to collectively denote all incoming and outgoing momenta, and $s_{i}$ as the signs in front of each term in Eq. (55) (+1+1 for the incoming momenta, 1-1 for the outgoing ones). The leading term in the soft expansion of the source is then
@@
which is consistent with Sachs’s peeling property (3). [^16] The appearance of the vector $L$ in the denominator is a consequence of the expansion at large $|\vec{x}|$. Indeed, writing $x^{\mu}=t\,t^{\mu}+|\vec{x}|r^{\mu}=ut^{\mu}+\sqrt{2}|\vec{x}|L^{\mu}$, the natural denominator becomes $(Q_{i}\cdot x)^{2}-Q_{i}^{2}x^{2}=|\vec{x}|^{2}(L\cdot Q_{i})^{2}(1+{\cal O}(u/|\vec{x}|)$. The appearance of higher powers of $1/|\vec{x}|$ in $\Psi_{0}$ and $\Psi_{1}$ may appear mysterious. For this contribution, it is a consequence of momentum conservation and projection onto the null tetrad (4), as we have derived this by making explicit use of the simple form of the leading terms in the soft expansion. In general, however, the relevant integrals are too difficult (and also not necessary) to evaluate. We will systematically understand the suppression of the $\Psi_{1}$ and $\Psi_{0}$ scalars with respect to the other Newman-Penrose scalars in Sec. 4.3, as a consequence of the angular integration.

### v2

@@ 3 The asymptotic expansion of expectation values
While Eq. (21) defines the metric sourced by a scattering process at all space-time points, carrying out the integrals can be difficult. For our purpose, however, such a calculation is unnecessary. Indeed, a large-$|\vec{x}|$ asymptotic expansion of this metric should be sufficient to probe its peeling properties. Furthermore, it is more convenient to evaluate Newman-Penrose scalars to probe gauge-invariant properties of spacetime. In this section, we begin by discussing the asymptotic expansion of Eq. (21) large values of $|\vec{x}|$. Then, we upgrade the analysis to cover the evaluation of the asymptotic expansion of derivative operators acting on the metric and, in particular, of the Newman-Penrose scalars.
@@
$k^{\mu}=\omega\,t^{\mu}+\mathsf{k}\,n^{\mu}\ ,\quad\text{where}\ n^{2}=-1\text{ and }t\cdot n=0$ (28)
@@
$\begin{split}\int\!\frac{d^{d}k}{(2\pi)^{d}}\frac{e^{-ik\cdot x}\,{\widetilde{h}}^{\mu\nu}(k)}{(k_{0}+i\epsilon)^{2}-\vec{k}^{2}}=\int_{-\infty}^{+\infty}\!\frac{d\omega}{2\pi}\int_{0}^{\infty}\!\frac{d\mathsf{k}}{2\pi}\frac{e^{-i\omega t\cdot x}\mathsf{k}^{d-2}}{(\omega+i\epsilon)^{2}-\mathsf{k}^{2}}\int\!dn\,e^{-i\mathsf{k}n\cdot x}{\widetilde{h}}^{\mu\nu}(k)\ ,\end{split}$ (29)
@@
The contour for the $\omega$ integral, as defined by the $i\epsilon$ prescription in Eq. (21), encodes our desired (retarded) boundary conditions. The integration can be carried out via a contour deformation. In particular, since we are mainly focused on contributions in the far future, by deforming the $\omega$ integration contour to the lower-half plane, we pick up contributions from the singularities of the source and the external (retarded) propagator. This integral can typically be performed exactly.
@@
This additional direction provides a natural parametrization for the graviton momentum:
@@
where $k^{\mu}$ in the argument of the current is given by Eq. (31), and $dn_{\perp}$ is the measure on the corresponding (d3)(d-3)-dimensional sphere.
The $\mathsf{k}$ integration is more subtle. At large distances, the integrand depends on two parametrically-separated scales: that of the system, $R_{i}\lesssim b$, and that of the observer, $|\vec{x}|$. Depending on the scaling of $\mathsf{k}$ with the observer scale, the contribution of the oscillatory factor is different. Thus, we have two distinct integration regions,
@@
Since the complete $\mathsf{k}$ integral is convergent both at small $\mathsf{k}$ (because of the known soft behavior of ${\widetilde{h}}^{\mu\nu}(k)$) and at large $\mathsf{k}$ (because of the exponential suppression ensuing from the typical frequency of ${\widetilde{h}}^{\mu\nu}(k)$ being ${\cal O}(1/b)$), a divergence in one of the two regions should be regularized by the cutoff $\lambda$ and must cancel against a similar divergence in the other region. Both of these regions contribute to the asymptotic expansion at large $|\vec{x}|$ and need to be considered separately. From the perspective of the method of regions in dimensional regularization Beneke:1997zp ; Jantzen:2011nz applied on the LHS of (29), the graviton momentum scales at large $|\vec{x}|$ in these regions are
$\text{Radiation region}:k^{\mu}\sim|\vec{x}|^{0}\ ,\qquad\qquad\text{Coulomb region}:k^{\mu}\sim|\vec{x}|^{-1}\ .$ (34)
@@ 3.1 The radiation region
In this region, it is convenient to deform the $z$ contour into two contours parallel to the imaginary axis and connected at infinity, as in Fig. 2. The integral takes the following form:
@@
which is valid for a vector $n$ parametrizing the unit $n$-sphere. In our case, $n_{\perp}$ is further constrained to be orthogonal on $t^{\mu}$ and $r^{\mu}$; the relevant projector enforcing this transversality is
@@
and the integral is a function of only $a^{\mu}_{\perp}=\eta_{\perp}^{\mu\nu}a_{\nu}$. The integrals over $\mathsf{k}$ and $\omega$ are more interesting. They depend on the details of the source ${\widetilde{h}}^{\mu\nu}(k)\sim J^{\mu\nu}(k)$ and are evaluated on a case-by-case basis. At tree-level, the $\omega$ integral can be evaluated via the residue theorem, as the only relevant singularities in the lower-half plane are the one in the external (retarded) propagator. Beyond tree level, the integral over $\omega$ is complicated by the appearance of logarithmic branch-points at $\omega=\mathsf{k}$ in $J^{\mu\nu}(k)$ in $d=4$. We leave the evaluation of these integrals to Sections 4.3 and 4.4. As a cross-check, we will also show there that the leading contribution in $|\vec{x}|^{-1}$ reproduces the waveform at future null infinity, originally obtained in Ref. Cristofoli:2021vyo via a saddle-point approximation.
@@ 3.2 The Coulomb region
In this region, the external graviton momentum is taken to be parametrically soft, $\sim|\vec{x}|^{-1}$; moreover, since we are interested in the metric at large distance towards future null infinity, we will assume that, while $\omega\neq\mathsf{k}$, they are of the same order, $\omega\sim\mathsf{k}$. Under these assumptions the source ${\widetilde{h}}^{\mu\nu}(k)$ can be expanded in powers of $\omega$ and $\mathsf{k}$. Displaying only the $\mathsf{k}$ dependence, ${\widetilde{h}}^{\mu\nu}(k)$ schematically has the general form Georgoudis:2023eke
@@
Since $\omega,\mathsf{k}\sim|\vec{x}|^{-1}$ in the Coulomb region, the phase of the Fourier transform in Eq. (3.1) must be expanded in the small quantity $\omega u$:
@@
Subleading orders contribute beyond the leading order in the $1/|\vec{x}|$ expansion. Following the usual procedure provided by the method of regions in dimensional regularisation, we take the upper limit of the integration domain to $\lambda\rightarrow\infty$ after series expanding the integrand.
@@
In the following, we will focus only on the leading and subleading soft contributions in Eq. (38). We will study both terms at all orders in the coupling. We find that only the subleading order in the soft expansion contribute to peeling violation. In particular, we will see that the leading $\mathsf{k}^{-1}$ contribution satisfies non-trivially the peeling conditions (3) and that the leading term in the ultra-soft exchange region reproduces at leading order in $G$ the peeling violation (114), identified in Ref. Christodoulou:2002 .
@@
are chosen so that the integrals in $J^{\mu\nu}(q_{1},q_{2})$ exhibit a clean separation of the classical orders. $J^{\mu\nu}(q_{1},q_{2})$ is the final-state graviton one-point function in momentum space, and the two regions correspond to the following rescalings Georgoudis:2023eke :
$\text{SE}:~(k^{\mu},q_{i}^{\mu})\sim(|\vec{x}|^{-1},1)\ ,\qquad\qquad\text{U-SE}:~(k^{\mu},q_{i}^{\mu})\sim(|\vec{x}|^{-1},|\vec{x}|^{-1})\ .$ (42)
SE and U-SE stand for soft-exchange and ultra-soft-exchange, respectively. As usual in the method of regions in dimensional regularization, the integrals are evaluated by expanding at small $\omega$ in each region and integrating over the entire integration domain.
@@ 3.3 Ansatz for metric
As discussed, gauge invariance of the Einstein-Hilbert action coupled to the source $J^{\mu\nu}$ requires the source to be transverse. We thus parametrize it in terms of manifestly-transverse tensors. At lowest order in the graviton momentum, there are six such tensors,
@@
The coefficients $\alpha_{i}$ turn out to be finite throughout the radiation region, $k^{\mu}\sim|\vec{x}|^{0}$, a property that we will leverage in later sections. To see this, recall that singularities of amplitudes (and of Green’s functions) correspond to physical processes and, as such, can be understood using a combination of unitarity arguments and analytic continuation Coleman:1965xm . In two-body scattering, the relevant invariants appearing in the impact-parameter-space form of the amplitude are $u_{1,2}\cdot k$, $b\cdot k$, $b^{2}$, $u_{1}\cdot u_{2}$, and $k^{2}$ (with the latter vanishing on-shell). The large distance expansion in the radiation region, corresponds to an essentially null momentum $k^{\mu}$, aligned with the null direction connecting the scattering event to the observer. In this limit, no singularities can arise. Indeed, the $u_{i}^{\mu}$ are time-like so divergences in $u_{i}\cdot k$ reduce to the standard soft singularities in $\omega$. Furthermore, no singularities can occur in $k\cdot b$, as the observer may lie anywhere on the celestial sphere and the impact parameter remains arbitrary. Moreover, since none of the tensor structures in Eq. (45) vanish as $z\rightarrow 1$, it follows that the coefficient functions $\alpha_{i}$ remain finite in this limit. These expectations can be confirmed explicitly from the known LO and NLO post-Minkowskian scattering waveforms Kovacs:1977uw ; Jakobsen:2021smu ; Herderschee:2023fxh ; Brandhuber:2023hhy ; DeAngelis:2023lvf ; Georgoudis:2023lgf ; Bini:2023fiz ; Brunello:2024ibk .
@@ 4 The peeling property by region
Using the methods outlined in the previous sections, we explicitly evaluate the asymptotic expansion of the Fourier transform (29). We discuss separately the Coulomb and radiation regions, focusing in Secs. 4.1, 4.2.1 and 4.3 on terms which are at rational in $\mathsf{k}$ and $\omega$. These originate from the soft-exchange part (42) of the Coulomb region (34), and from the terms which are analytic in $k^{2}$ in the radiation region (34). We will prove that such terms satisfy Sachs’s peeling property (3) as a consequence of nontrivial cancellations coming from the angular integration. We then discuss the remaining contributions – the ultra-soft-exchange part (42) of the Coulomb region (34) in Sec. 4.2.2 and the non-analytic terms in $k^{2}$ in the radiation region (34) in Sec. 4.4. The former yields violation of the peeling property already identified at leading order in Newton’s constantin Refs. Christodoulou:2002 . The latter, which we discuss in Sec. 4.4, introduce a new stronger violation.
@@ 4.1 Asymptotic behavior of the waveform: the gravitational memory
is on the one hand subleading in the soft limit and, on the other, it is analytic. These properties allow us to ignore the difference at leading order. Direct power counting indicated that the off-shell deformation contributes only suppressed terms in the large-$|\vec{x}|$ expansion; moreover, as we will see, the fact that they are proportional to the off-shellness $k^{2}$ will also allow us to ignore them for the study of peeling properties. The leading soft behavior of the amplitude captures the information about asymptotic large retarded times $|u|$, i.e. the so-called memory effect.
The soft theorem for the two-body-scattering in-in correlator is [^16] Such a soft theorem for the two-body scattering can be inferred e.g from the result of Ref. Saha:2019tub : the leading soft theorem is given by $J^{\mu\nu}(k)=i\frac{\kappa^{2}}{2}\sum_{i=1}^{n+m}\frac{p_{i}^{\mu}p_{i}^{\nu}}{p_{i}\cdot k+i\epsilon}\ ,$ (53) where $n$ and $m$ are the number of incoming and outgoing particles, respectively (either massive or massless; for the incoming particles the $p_{i}$ are intended as minus their momenta). As our focus on the two-body scattering, we have to sum over the momenta of the incoming and outgoing massive states. During the scattering process, a certain amount of radiation is emitted and its total momentum is fixed by momentum conservation. Requiring gauge invariance to hold in the soft limit, fixes the fifth term to be of the form in Eq. (4.1), where $\rho(K)$ is the energy spectrum, given by modulus squared of the waveform (see Ref. Georgoudis:2025vkk for a recent discussion). Note also that we work in a mostly-minus metric convention while Ref. Saha:2019tub has the opposite, mostly-plus metric convention.
$J^{\mu\nu}(k)$ $=i\frac{\kappa^{2}}{2}\left[-\frac{p_{1}^{\mu}p_{1}^{\nu}}{p_{1}\cdot k-i\epsilon}-\frac{p_{2}^{\mu}p_{2}^{\nu}}{p_{2}\cdot k-i\epsilon}+\frac{p_{1}^{\prime\mu}p_{1}^{\prime\nu}}{p_{1}^{\prime}\cdot k+i\epsilon}+\frac{p_{2}^{\prime\mu}p_{2}^{\prime\nu}}{p_{2}^{\prime}\cdot k+i\epsilon}\right$.
$\qquad\quad\;\left.+\!\int\!d^{d}K\rho(K)\frac{K^{\mu}K^{\nu}}{K\cdot k+i\epsilon}\right]+\mathcal{O}(\mathsf{k}^{0},\mathsf{k}^{d-4})\ ,$ (54)
where $p_{1,2}^{\mu}$ are the two incoming matter momenta, $p_{1,2}^{\prime\mu}$ are (exact) outgoing momenta, and $K^{\mu}$ is the four-momentum carried by the emitted radiation during the scattering process. The outgoing matter momenta are determined by the details of the interactions between the two bodies, and we will not need their precise form, which is known only perturbatively. For example, at leading order in Newton’s constant
@@
$p_{1}^{\mu}+p_{2}^{\mu}=p_{1}^{\prime\mu}+p_{2}^{\prime\mu}+\!\int\!d^{d}K\rho(K)K^{\mu}\ .$ (56)
It is convenient to introduce the shorthand notation $Q_{i}^{\mu}$’s to collectively denote all incoming and outgoing momenta, and $s_{i}$ as the signs in front of each term in Eq. (4.1) (+1+1 for the incoming momenta, 1-1 for the outgoing ones). The leading term in the soft expansion of the source is then
@@
which is consistent with Sachs’s peeling property (3). [^19] The appearance of the vector $L$ in the denominator is a consequence of the expansion at large $|\vec{x}|$. Indeed, writing $x^{\mu}=t\,t^{\mu}+|\vec{x}|r^{\mu}=ut^{\mu}+\sqrt{2}|\vec{x}|L^{\mu}$, the natural denominator becomes $(Q_{i}\cdot x)^{2}-Q_{i}^{2}x^{2}=|\vec{x}|^{2}(L\cdot Q_{i})^{2}(1+{\cal O}(u/|\vec{x}|)$. The appearance of higher powers of $1/|\vec{x}|$ in $\Psi_{0}$ and $\Psi_{1}$ may appear mysterious. For this contribution, it is a consequence of momentum conservation and projection onto the null tetrad (4), as we have derived this by making explicit use of the simple form of the leading terms in the soft expansion. We will systematically understand the suppression of the $\Psi_{1}$ and $\Psi_{0}$ scalars with respect to the other Newman-Penrose scalars in Sec. 4.3, as a consequence of the angular integration.